OpenStudy (anonymous):

The length of a rectangle is 5 more than 3 times its width. The perimeter of the rectangle is 58 inches. What is the length of the rectangle? (Points : 4) 23 in. 29 in. 32 in. 35 in.

OpenStudy (zpupster):

2w + 2L = perimeter so perimeter is 58 the given length is 5+ 3w plug that in the formula 2w + 2(3w+5) = 58 2w + 6w + 10 = 58 we mult thru parentheses 8w = 48 we combined the the w terms and subtracted 10 from both sides divide 8 both sides that will give you the width plug it back in the 2w + 2L = perimeter to find the length if you need help ask

OpenStudy (anonymous):

23 in Solve: 2 (w + 3 w + 5) = 58 w = 6 Length = 3w + 5 = 18 + 5 = 23

OpenStudy (anonymous):

Thanks, do you think you could help me with one more? I am still having trouble.

OpenStudy (anonymous):

Sorry, it's 0205 Pacific Time. Have to bail out. Been on this site off and on all day.

OpenStudy (anonymous):

Thanks anyway!

OpenStudy (zpupster):

just post it

OpenStudy (anonymous):

The perimeter of a rectangular photo is 72 cm. The length is twice the width. What are the length and width of the photo?